OpenStudy (anonymous):

solve: 3sec^2(x)-4=0

OpenStudy (anonymous):

sec^2(x)=4/3 or sec (x)=2/(3)^(1/2) or x= pi/6

OpenStudy (dumbcow):

isolate the sec^2 \[\sec^{2} x = \frac{4}{3}\] take sqrt \[\sec x = \frac{2}{\sqrt{3}}\] remember that sec = 1/cos so by flipping both sides \[\cos x = \frac{\sqrt{3}}{2}\] find angle using unit circle or calculator

OpenStudy (anonymous):

ok. thanks @dumbcow ! :)

OpenStudy (dumbcow):

i should add there are 2 answers, i forgot the +- also find \[\cos x = -\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

ok